OpenStudy (anonymous):

marty leaves the airport in his private plane and flies due east at 186 mph. two hrs later a jet leaves the airport and flies due east at 434mph. when will the jet overtake marty's plane?

OpenStudy (anonymous):

no

OpenStudy (lilg132):

would you like the working out?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

in 1 hour

OpenStudy (anonymous):

i got 1.5

OpenStudy (anonymous):

x - hours after jet left \[s=vt\] \[186\cdot2+186\cdot x=434\cdot x\] \[x=1.5\]